How do you write an equation of a line using
Contains (5,8) and that is perpendicular to the line
Containing (0,-6) and (1,1)

Answer by mananth(16949)

x1 y1 x2 y2
0 -6 1 1

slope m = (y2-y1)/(x2-x1)
( 1 - -6 )/( 1 - 0 )
( 7 / 1 )
m= 7
slope of line perpendicular will be negative reciprocal of 7
-1/7
m= - 1/ 7

Plug value of the slope and point ( 5 , 8 ) in
Y = m x + b
8.00 = - 5/7 + b
b= 8 - - 5/7
b= 8 5/ 7
So the equation will be
Y = - 1/7 x + 61/7
m.ananth@hotmail.ca